OpenStudy (anonymous):

You want to pick a four-player tennis team. There are eight player to choose from. How many different teams can you form? A) 70 B) 1680 C) 106 D) 1210

OpenStudy (anonymous):

use combination

OpenStudy (anonymous):

8 C 4 = (8 * 7 * 6 * 5 )/(4 * 3 * 2 * 1) = 7 * 6 * 5 /3 = 7 * 2 * 5 = 70 \(\huge\cal\rlap{\color{lime}-\color{lime}G\color{lime}r\color{lime}e\color{lime}e\color{lime }n\color{lime}ツ}{\color{black}{\;-Green~ツ}}\)

OpenStudy (anonymous):

@msawesome111

OpenStudy (anonymous):

Thankk you so much!!

OpenStudy (anonymous):

Thank*

OpenStudy (anonymous):

No problem :)